### v2

# Enhancing Multimodal Medical Image Classification using Cross-Graph Modal Contrastive Learning

Abstract
The classification of medical images is a pivotal aspect of disease diagnosis, often enhanced by deep learning techniques. However, traditional approaches typically focus on unimodal medical image data, neglecting the integration of diverse non-image patient data. This paper proposes a novel Cross-Graph Modal Contrastive Learning (CGMCL) framework for multimodal medical image classification. The model effectively integrates both image and non-image data by constructing cross-modality graphs and leveraging contrastive learning to align multimodal features in a shared latent space. An inter-modality feature scaling module further optimizes the representation learning process by reducing the gap between heterogeneous modalities. The proposed approach is evaluated on two datasets: a Parkinson’s disease (PD) dataset and a public melanoma dataset. Results demonstrate that CGMCL outperforms conventional unimodal methods in accuracy, interpretability, and early disease prediction. Additionally, the method shows superior performance in multi-class melanoma classification. The CGMCL framework provides valuable insights into medical image classification while offering improved disease interpretability and predictive capabilities.
Index Terms:
Neurodegenerative disease, SPECT, Contrastive learning, Multimodal fusion, Classification, Cross-graph modal graph learning.
1 Introduction
In recent years, medical computing has increasingly shifted toward utilizing deep learning frameworks as a foundation for diagnostic analysis. Single convolutional neural networks (CNNs) have demonstrated substantial success in various unimodal medical imaging applications, such as computer-aided detection and diagnosis, image segmentation, and survival analysis [1, 2, 3]. However, the landscape of medical diagnostics has become increasingly complex with the growing availability of diverse data modalities. These include unstructured electronic health records (EHRs), quantitative blood test results, and demographic data, making it difficult to rely solely on single-modality data (e.g., clinical notes) for accurate disease predictions. [4, 5].
Injecting radioactive tracers for single-photon emission computed tomography (SPECT) scans remains a critical diagnostic tool for neurodegenerative diseases like Parkinson’s disease (PD), as it provides insights into the physiological functions and metabolic activities of organs beyond their anatomical structures [6]. However, SPECT has limitations, including relatively low spatial resolution and higher image noise due to the limited dose of radioactive tracers. Recent research has employed CNN-based models to address these issues, showing high accuracy in classifying and predicting PD [7, 8, 6]. Despite their success, these models often fail to incorporate or provide insights into relevant patient demographics in clinical practice.
Graph-based approaches offer a promising avenue for identifying commonalities among patients’ medical imaging data and symptom profiles [9, 10, 11]. Specifically, graph convolutional neural networks (GCNs) [12] have demonstrated remarkable efficacy in processing the complex interconnections and structural patterns inherent in non-linear data structures like graphs [11, 13]. While traditional machine learning methods often struggle with such sophisticated data, GCNs excel at interpreting and analyzing these relationships. GCNs learn from unstructured features by processing non-Euclidean distances, enabling them to perform various downstream tasks, including node classification, link prediction, and graph classification. An advanced framework of GCNs includes graph attention networks (GATs) [14], which introduce an attention mechanism to enhance the learning and fusion of neighboring node features. GATs can capture correlations between modalities, establish adaptive graph learning from heterogeneous multimodal feature spaces, and ultimately perform disease prediction in downstream tasks [15].
Contrastive learning, a self-supervised framework, aims to minimize the distance between representations of positive sample pairs while maximizing it for negative pairs. This approach offers a promising solution to handling multimodal data similarity, complicated by the heterogeneous structural characteristics across different modalities [16, 17]. This approach aligns multimodal data in a shared embedding space for Alzheimer’s disease (AD) prediction. Additionally, attention mechanisms applied to tabular data improve model performance and interpretability [18], while multimodal transformers that combine image and clinical data can predict AD progression. [19]. However, these studies have typically focused on a specific disease, with limited research on model generalization across different multimodal disease datasets. This study proposes a novel approach that leverages dual graph networks to construct a cross-graph modal feature topology for diverse modalities, such as medical images and quantitative parameters. Our method employs GATs to learn feature encodings across modalities through node embedding. Additionally, we design a graph contrastive loss to enhance the similarity of the final representations. We validate our proposed Cross-Graph Modal Contrastive Learning (CGMCL) approach for multimodal fusion as shown in Fig. 2, and clinical interpretability using two multimodal medical datasets: a private PD dataset and a public melanoma dataset.
Figure 1: The four neural network multimodal fusion methods are as follows: (a) and (b) represent conventional and widely-used vectors, with (a) utilizing vector concatenation and (b) employing attention-based modal learning. Method (c) uses a joint network for feature extraction from diverse modalities. Finally, method (d) illustrates our proposed cross-graph modal fusion, incorporating a graph structure.
Figure 2: The framework of multimodal cross-graph fusion for constructing a common feature space with contrastive learning.
2 RELATED WORK
2.1 Multimodal Learning in Medical Imaging
Image-based models have traditionally relied on unimodal input, primarily for disease classification. However, structured medical images alone are often insufficient for integrating patients’ physiological or numerical characteristics. Recent research has demonstrated the effectiveness of multimodal deep-learning approaches. For instance, a multi-co-attention model successfully integrated brain SPECT images with DNA methylation data [20]. In multimodal survival analysis, hazard functions are estimated by integrating features from co-attention modules applied to diverse pathology images and genetic data [2, 21]. However, in the above studies, feature extraction from medical images has predominantly been conducted at the pixel level, with limited consideration for structured feature learning that accounts for inter-patient image similarities (e.g., graph structures). Therefore, we propose that early diagnosis, particularly for conditions such as PD, requires more precise multimodal models to improve early prediction and clinical interpretability.
2.2 Multi-Modality Fusion Methods
The greatest challenge lies in effectively fusing modalities from different domains. Fig. 1 illustrates various feature fusion methods, including feature extraction from two modalities, followed by fusion approaches such as vector concatenation, attention-based fusion, and joint network fusion learning to combine cross-domain features [22], as shown in Fig. 1 (a)-(c). However, those fusion methods do not account for calculating similarities between different modalities. We first construct a graph structure using image feature vectors extracted from a pre-trained ResNet model as single-modality input features for each graph node to address this. For example, a GCN can use patch images from areas of pathological interest after gridding, with nodes representing patients and edges representing correlations between them [23, 24]. In this study, we propose a cross-graph model that inputs features from two separate encoder modalities. Subsequently, we construct a common modality embedding to better fuse the two domain latent spaces, as illustrated in Fig. 1 (d).
3 METHODOLOGY
3.1 Problem Definition
This study focuses on two medical modalities: medical images and non-image data (e.g., meta-features). We denote a multimodal set $\textbf{X}=\{X_{1},X_{2},...,X_{N}\}$ for $N$ patients. We can represent the data for the $i$-th patient as a 3-tuple $X_{i}=\{(I_{i},C_{i},Y_{i})\}_{i=1}^{N}$, where $I_{i}\in\mathbb{R}^{h\times h}$ represents the medical image with a scale dimension $h$, $C_{i}\in\mathbb{R}^{F}$ represents the meta-features with $F$ features, and $Y_{i}$ corresponds to the disease labels. We first construct a non-linear model (e.g., an CNN) to generate initial feature maps $I_{i}^{\prime}$ from the images:
$I_{i}^{\prime}=f(\cdot)=\mathrm{CNN}(I_{i}),$ (1)
where the function $f(\cdot):\mathbb{R}^{D}\to\mathbb{R}^{d}$ represents the feature extractor for images, $D$ is the dimension of the initial latent space in the $l-1$ layer, and $d$ is the reduced feature dimension after the $l$-th convolution layer.
3.2 Graph Construction
In practice, to better measure the structural properties of the two modalities in non-Euclidean distance feature spaces, we consider the two modality graphs $\mathcal{G}^{I}(\mathcal{E}^{I},\mathcal{V}^{I})$ and $\mathcal{G}^{C}(\mathcal{E}^{C},\mathcal{V}^{C})$, with edges $\left|\mathcal{E}^{I}\right|$ and $\left|\mathcal{E}^{C}\right|$ and vertices $\left|\mathcal{V}^{I}\right|$ and $\left|\mathcal{V}^{C}\right|$, respectively. Given the input encoded features $F_{i}$ (e.g., $I_{i}^{\prime}$ or $C_{i}$), we can construct the binary adjacency matrices $A^{I}$ and $A^{C}$ using a $K$-nearest neighbors graph [25]:
[…]
where $\mathcal{N}(\cdot)$ denotes the set of indices of the $K$ nearest neighbors of features $F_{i}$ and $F_{j}$ based on Euclidean distance. In a $K$-neighborhood, two data points $i$ and $j$ are connected by an edge $\mathcal{E}_{(i,j)}$ if $i$ is among the $K$ nearest neighbors of $j$, or vice versa. Each vertex $\mathcal{V}$ in the graph represents a patient.
3.3 Graph Attention Encoder
Our modality-based graph model employs a GAT as the foundational features encoder to facilitate cross-graph modal learning across different graph modalities. Given the two input features node representations $I^{\prime}=[I^{\prime}_{1},I^{\prime}_{2},...,I^{\prime}_{N}]\in\mathbb{R}^{N\times F^{\prime}}$ and $C=[C_{1},C_{2},...,C_{N}]\in\mathbb{R}^{N\times F^{\prime}}$, we can formulate the multi-modality attention coefficients in a GNN as:
$e^{I^{\prime}}_{ij}=\sigma\left(\mathbf{W}^{I^{\prime}}I^{\prime}_{i},\mathbf{W}^{I^{\prime}}I^{\prime}_{j}\right),$ (3)
$e^{C}_{ij}=\sigma\left(\mathbf{W}^{C}C_{i},\mathbf{W}^{C}C_{j}\right),$ (4)
where $\sigma(\cdot)$ represents a non-linear transformation function (e.g., LeakyReLU, tanh), and the trainable weighted matrices are $\textbf{W}^{I^{\prime}}$, $\textbf{W}^{C}$. We can then normalize the attention coefficients across neighboring nodes using the softmax function, which can be expressed as:
$\alpha^{I^{\prime}}_{ij}=\mathrm{softmax}_{j}(e^{I^{\prime}}_{ij})=\frac{\exp(e^{I^{\prime}}_{ij})}{\sum_{k\in\mathcal{N}_{i}}\exp(e^{I^{\prime}}_{ik})},$ (5)
$\alpha^{C}_{ij}=\mathrm{softmax}_{j}(e^{C}_{ij})=\frac{\exp(e^{C}_{ij})}{\sum_{k\in\mathcal{N}_{i}}\exp(e^{C}_{ik})},$ (6)
We apply weighted aggregation to the neighborhood node vectors using the normalized attention coefficients as attention scores:
$\vspace{-1mm}H_{i}^{I^{\prime}}=\sigma\left(\sum_{j\in\mathcal{N}_{i}}\alpha^{I^{\prime}}_{ij}\mathbf{W}^{I^{\prime}}{I^{\prime}}_{j}\right),$ (7)
$H_{i}^{C}=\sigma\left(\sum_{j\in\mathcal{N}_{i}}\alpha^{C}_{ij}\mathbf{W}^{C}C_{j}\right),$ (8)
where $H^{I^{\prime}}$ and $H^{C}$ are the output representations from the GAT encoder.
We focus on enhancing the fusion of learned representations across cross-graph node modality embeddings. To integrate both the GAT output representations $H_{i}^{I^{\prime}}$ and $H_{i}^{C}$ with the encoder feature information, we also incorporate a concatenation of features generated by the CNN encoder $I^{\prime}$ and the meta-features $C$ into the graph encoder. This can be expressed as:
$\mathcal{H}^{I}=\left[H_{i}^{I}\mathbin{\|}I_{i}^{\prime}\right],$ (9)
$\mathcal{H}^{C}=\left[H_{i}^{C}\mathbin{\|}C_{i}\right],$ (10)
[…]
3.4 Inter-Modality Feature Enhancement and Scaling
To address scale discrepancies between features and reduce modality gaps in the latent space for two single-modality output features, we introduce the inter-modality feature enhancement and scaling (IMFES) module. This module is designed to enhance the learning of intrinsic modality distributions and preserve important structural information within each modality. The process begins by applying a simple multilayer perceptron (MLP) to the concatenated tensors $\mathcal{H}^{I}$ and $\mathcal{H}^{C}$, transforming them into a single-modality probability matrix. Next, we perform element-wise multiplication between the GAT encoder outputs $H^{I^{\prime}}$ and $H^{C}$, and the non-linear MLP transforms $\zeta(\mathcal{H}^{I})$ and $\zeta(\mathcal{C}^{I})$, where $\zeta(\cdot)=\mathrm{MLP(\cdot)}$. This operation enables fine-grained extraction of individual modality features, as represented by the following equations:
$E^{I}=H^{I^{\prime}}\odot\zeta(\mathcal{H}^{I}),$ (11)
$E^{C}=H^{C}\odot\zeta(\mathcal{H}^{C}),$ (12)
where $E^{I}$ and $E^{C}$ are the resulting scaled feature matrices for the respective modalities, enhancing the representations learned by the GAT encoders, and $\odot$ denotes element-wise multiplication. This operation mitigates the effects of disparities in scale or distribution shape between the two sets of features.
3.5 Cross-Graph Modal Fusion
To integrate features from cross-graph modality representations at varying scales, we concatenate each image embedding $E^{I}$ with its corresponding representation $\mathcal{H}^{I}$ to form $\hat{Z}^{I}=[\mathcal{H}^{I}\mathbin{\|}E^{I}]$, and similarly for clinical data, $\hat{Z}^{C}=[\mathcal{H}^{C}\mathbin{\|}E^{C}]$, resulting in final embedding matrices. We then combine these cross-modal embeddings by element-wise addition and apply a sigmoid function: $\hat{Z}=\sigma(\hat{Z}^{I}+\hat{Z}^{C})$. This process aligns representations from both modalities in a common feature space. Furthermore, to integrate the feature spaces of images and meta-features in the same embedded space, we construct a similarity matrix $\textbf{S}\in\mathbb{R}^{N\times N}$ for each pair of similar patients using the final embedding $\hat{Z}$ learned from the model. The similarity between the $i$-th and $j$-th patients can be defined as follows:
$S_{ij}=\hat{Z}_{i}\cdot(\hat{Z}_{j})^{T},\forall i,j\in\left[1,N\right],$ (13)
where $S_{ij}$ represents the similarity between patients $i$ and $j$, incorporating information from both modalities.
3.6 Contrastive Loss Optimization
[…]
$\Theta(a)=\begin{cases}1,&\text{if }a_{ij}\geq 0,\\ 0,&\text{if }a_{ij}<0,\end{cases}$ (14)
where $a_{ij}$ represents elements in $\hat{A}^{I}$ or $\hat{A}^{C}$.
We can calculate the positive and negative pairs in the similarity matrix as $S^{+}=S\odot D_{pos},\quad S^{-}=S\odot D_{neg}$. The sum of the positive and negative scores can then be calculate as:
$P_{s}=\sum_{i=1}^{N}\sum_{j=1}^{N}S_{ij}^{+}Y_{j1},$ (15)
[…]
where $\delta>0$ is the controllable margin, and $Y$ is the label matrix. The positive loss and negative loss can then be written as:
$\mathcal{L}_{pos}=-\sum_{i=1}^{N}\log\left(P_{s}+\epsilon\right),$ (17)
$\mathcal{L}_{neg}=-\sum_{i=1}^{N}\log\left(N_{s}+\epsilon\right),$ (18)
where $\epsilon$ is $1\times 10^{-8}$ is used to prevent numerical computation issues. By using Eq. 17 and 18 we can ultimately obtain the combined losses, incorporating both the positive and negative loss, written as $\mathcal{L}_{contrastive}=\mathcal{L}_{pos}+\mathcal{L}_{neg}$. By minimizing $\mathcal{L}_{contrastive}$, the intra-class similarity is maximized, and the inter-class dissimilarity is increased.
To optimize the loss function and predict the probabilities of the final disease classes, we incorporated both $\hat{Z}^{I}$ and $\hat{Z}^{C}$ into the supervised binary classification loss using the softmax function. The cross-entropy loss function can be expressed as:
$\mathcal{L}_{I}=-\sum_{i=1}^{N}y_{i}^{T}\ln(\text{softmax}(\hat{y}_{i}^{I})),$ (19)
$\mathcal{L}_{C}=-\sum_{i=1}^{N}y_{i}^{T}\ln(\text{softmax}(\hat{y}_{i}^{C})),$ (20)
where $y_{i}$ is the one-hot vector of the true label, and $\hat{y}_{i}^{I}$ and $\hat{y}_{i}^{C}$ are the model’s outputs for the image and meta-features, respectively. During the optimization process, we developed a comprehensive loss function that integrates cross-entropy and contrastive loss from the two cross-graph modalities. To further enhance the effectiveness of this combined loss, we incorporated the mean squared error between the similarity matrix $S$ and the diagonal matrix $D_{ii}=\sum_{i}A_{ii}$ when calculating the clustering of the module output fusion. The extend diagonal loss is expressed as:
$\mathcal{L}_{diag}=\frac{1}{N}\sum^{N}_{i,j}(S_{ij}-D_{ii})^{2},$ (21)
We use $\beta$ as a leverage coefficient to control the optimization weight of the overall loss, which is defined as follows:
$\mathcal{L}_{CGMCL}=(1-\beta)(\mathcal{L}_{m}+\mathcal{L}_{f})+\beta\mathcal{L}_{contrastive}+\mathcal{L}_{diag},$ (22)
where $\beta$ can be set between 0 and 1, and it controls the contribution level of different losses. The coefficient $\beta$ adjusts the weight assigned to each loss component.
TABLE I: Summary statistics of demographics in PD patients with different subtypes
Subtypes No. of Subjects Male(%) Age (Mean $\pm$ Std)
Normal / Abnormal 127 / 154 54.6% 67.5 $\pm$ 11.2
Normal / MA 127 / 131 44.1% 68.0 $\pm$ 12.0
MA / Abnromal 131 / 154 55.8% 68.4 $\pm$ 11.3
4 Dataset Collection
  • Parkinson’s disease (PD) Data for this study was collected at Kaohsiung Chang Gung Memorial Hospital, Taiwan, from January 2017 to June 2019, involving 416 patients [26]. The study received approval from the Chang Gung Medical Foundation Institutional Review Board, and all data were de-identified. Four expert nuclear medicine physician annotated the data, labeling PD across three subtypes: Normal, Mildly Abnormal (MA), and Abnormal. Details regarding the annotation criteria can be found in Supplementary Fig. S2, while Table I summarizes the descriptive statistics. The images used in this study were Tc99m TRODAT single-photon emission computed tomography (SPECT) scans acquired using a hybrid SPECT/CT system (Symbia T, Siemens Medical Solution). Image acquisition involved 30-second steps across 120 projections, covering a full 360-degree circular rotation with low-energy, high-resolution parallel-hole collimators. After reconstruction, CT-based attenuation correction imported the images into DaTQUANT for automatic semi-quantification of the DaTQUANT meta-features [27]. Twelve parameters were obtained from DaTQUANT: Striatum Right (S-R), Striatum Left (S-L), Anterior Putamen Right (AP-R), Anterior Putamen Left (AP-L), Posterior Putamen Right (PP-R), Posterior Putamen Left (PP-L), Caudate Right (C-R), Caudate Left (C-L), Putamen/Caudate Ratio Right (P/C-R), Putamen/Caudate Ratio Left (P/C-L), Putamen Asymmetry (PA), and Caudate Asymmetry (CA). Supplementary Fig. S1 contains additional details about these meta-features. The original SPECT images (800 × 1132) were resized to a standardized resolution of 128 × 128 pixels for model development. A total of 412 preprocessed images and their twelve associated quantitative DaTQUANT meta-features were utilized for model training (n = 300) and testing (n = 112).
  • Melanoma dataset [28] The melanoma open dataset utilized for this study is a publicly available 7-point multimodal dataset comprising dermoscopic images and clinical data from 413 training and 395 testing samples. The classification task involves seven key image-based features: 1) pigment network (PN), 2) blue whitish veil (BWV), 3) vascular structures (VS), 4) pigmentation (PIG), 5) streaks (STR), 6) dots and globules (DaG), and 7) regression structures (RS). Additionally, the dataset includes five diagnostic categories: 1) basal cell carcinoma (BCC), 2) blue nevus (NEV), 3) melanoma (MEL), 4) miscellaneous (MISC), and 5) seborrheic keratosis (SK). The dermoscopic images have a resolution of 512 × 768 pixels, while the clinical data contains information on the patient’s gender and lesion location. More melanoma categories and their annotation information can be referenced in Table 1 of the literature  [28]
5 EXPERIMENTS
5.1 Baseline Methods Comparison
To fairly compare the effectiveness of CGMCL across various methods for both the PD and melanoma datasets, we conducted a quantitative analysis of baseline method comparisons. The results for PD, segmented into three subtypes, are presented in Table II through Table IV, and for the melanoma multi-class dataset in Table V. Initially, we employed conventional machine learning algorithms to assess the impact of a single modality on meta-features. This analysis applied logistic regression, XGBoost, random forest, support vector machines (SVM) [29], and AdaBoost [30] to both datasets. To further evaluate the performance of our proposed CGMCL model in PD classification, we compared it to unimodal two-layer CNN models. We then extended the analysis using SPECT images with three slices to compare different 3D-CNN architectures. When comparing machine learning methods that utilize only meta-features against CNN-based unimodal approaches, nonlinear logistic and tree-based methods (i.e., XGBoost and AdaBoost) achieved accuracies between 0.58 and 0.63 for distinguishing Normal from MA in PD. This indicates that relying solely on quantitative data may result in losing other informative features, such as those derived from images. In contrast, unimodal image feature extraction models substantially improved classification accuracy for the three PD subtypes. Specifically, ResNet18 increased accuracy by 16% for Normal vs. MA, 12.5% for Normal vs. Abnormal, and 1.16% for Normal vs. Abnormal compared to logistic regression. Conversely, for the melanoma dataset, meta-features in ML methods outperform in the DaG, PN, and VS categories.
Figure 3: The ablation study of the weighted loss coefficient $\beta$ is shown in Panel (a), while Panel (b) illustrates the convergence of the two modality features’ alignment using KL divergence.
5.2 Compare with the SOTA Methods
  1. 1.
    Compared to unimodal CNN-based feature extractors (i.e., 2-layer CNN, 3D-CNN, VGG19, and ResNet18), the three SOTA models using multimodal 3D-CNNs did not demonstrate substantially higher ACC across the three PD subtypes. However, individual metrics improved for certain models. Specifically, DeAF showed enhancements in specificity (SPE) and positive predictive value (PPV) for Normal vs. MA (SPE = 0.76) and for MA vs. Abnormal (SPE = 0.91, PPV = 0.93). MHCA also improved sensitivity (SEN) for Normal vs. Abnormal (SEN = 0.92).
  2. 2.
    We further evaluated the cross-modal graph fusion capability of our proposed CGMCL model in classifying the three PD subtypes. From Table II, using GAT as the cross-modal fusion method with ResNet18 achieves metric values between 0.73 and 0.76, comparable to the SOTA model using 3D-CNNs or unimodal approaches (ACC $\uparrow$ 4.1%, PPV $\uparrow$ 3.9%). This indicates that the fusion of clinical and imaging data can more effectively assist in identifying difficult-to-diagnose early PD symptoms. Moreover, there is a noticeable improvement in classification performance for the other two subtypes (e.g., MA vs. Abnormal and Normal vs. Abnormal).
  3. 3.
    When comparing our proposed CGMCL with other SOTA models on the melanoma dataset, although CGMCL’s two-class classification performance for BWV and RS was lower than that of the latest FusionM4Net model, it achieved higher accuracy in most three-class classifications (DaG, PIG, PN, VS, and DIAG). This superiority in multi-class tasks is due to the model’s ability to handle imbalanced sample issues and fine-grained categories across multiple classes. Additionally, our CGMCL model demonstrated consistent performance across multi-class challenges. Our approach achieved an impressive ACC of 0.95 for the five-class DIAG task.
TABLE II: The comparison of the proposed CGMCL’s performance (mean $\pm$ std) between unimodal and multimodal approaches for PD subtypes: Normal vs MA. The best performance results are highlighted in bold, and “_” indicates the second-best methods
Normal vs. MA
Models Backbone Image Feature Extractor Clinical Features ACC SEN SPE PPV NPV
Logistic - - 0.63 $\pm$ 0.00 0.65 $\pm$ 0.01 0.62 $\pm$ 0.01 0.66 $\pm$ 0.02 0.61 $\pm$ 0.03
XGboost - - 0.58 $\pm$ 0.00 0.56 $\pm$ 0.01 0.59 $\pm$ 0.02 0.61 $\pm$ 0.03 0.55 $\pm$ 0.03
AdaBoost - - 0.56 $\pm$ 0.07 0.58 $\pm$ 0.12 0.55 $\pm$ 0.08 0.57 $\pm$ 0.08 0.56 $\pm$ 0.08
Single slice (2-D)
- 2-layer CNN 0.65 $\pm$ 0.05 0.65 $\pm$ 0.06 0.64 $\pm$ 0.06 0.67 $\pm$ 0.06 0.62 $\pm$ 0.06
Unimoal - VGG19 0.71 $\pm$ 0.02 0.72 $\pm$ 0.03 0.70 $\pm$ 0.03 0.73 $\pm$ 0.03 0.69 $\pm$ 0.04
- ResNet18 0.73 $\pm$ 0.05 0.75 $\pm$ 0.06 0.69 $\pm$ 0.06 0.74 $\pm$ 0.05 0.72 $\pm$ 0.06
Three slices (3-D)
MHCA [31] - 3D-CNN 0.58 $\pm$ 0.02 0.52 $\pm$ 0.02 0.68 $\pm$ 0.03 0.68 $\pm$ 0.03 0.51 $\pm$ 0.02
DeAF [32] - 3D-CNN 0.62 $\pm$ 0.02 0.51 $\pm$ 0.11 0.76 $\pm$ 0.10 0.76 $\pm$ 0.06 0.54 $\pm$ 0.03
TriFormer [19] Transformer 3D-CNN 0.63 $\pm$ 0.02 0.56 $\pm$ 0.04 0.73 $\pm$ 0.02 0.74 $\pm$ 0.02 0.55 $\pm$ 0.02
Proposed Model (2-D)
CGMCL GCN+GCN ResNet18 0.70 $\pm$ 0.05 0.66 $\pm$ 0.06 0.75 $\pm$ 0.08 0.75 $\pm$ 0.08 0.66 $\pm$ 0.05
CGMCL GCN+GCN VGG19 0.65 $\pm$ 0.03 0.65 $\pm$ 0.05 0.65 $\pm$ 0.02 0.67 $\pm$ 0.03 0.63 $\pm$ 0.04
CGMCL GAT+GAT VGG19 0.70 $\pm$ 0.04 0.69 $\pm$ 0.06 0.69 $\pm$ 0.04 0.71 $\pm$ 0.05 0.66 $\pm$ 0.05
CGMCL GAT+GAT ResNet18 0.76 $\pm$ 0.03 0.75 $\pm$ 0.03 0.78 $\pm$ 0.05 0.79 $\pm$ 0.05 0.73 $\pm$ 0.04
TABLE III: The comparison of the proposed CGMCL’s performance (mean $\pm$ std) between unimodal and multimodal approaches for PD subtypes: Abnormal vs Abnormal. The best performance results are highlighted in bold, and “_” indicates the second-best methods
MA vs. Abnormal
Models Backbone Image Feature Extractor Clinical Features ACC SEN SPE PPV NPV
Logistic - - 0.86 $\pm$ 0.00 0.88 $\pm$ 0.02 0.84 $\pm$ 0.02 0.85 $\pm$ 0.01 0.87 $\pm$ 0.01
XGboost - - 0.83 $\pm$ 0.00 0.83 $\pm$ 0.00 0.82 $\pm$ 0.02 0.83 $\pm$ 0.02 0.82 $\pm$ 0.01
AdaBoost - - 0.75 $\pm$ 0.03 0.73 $\pm$ 0.09 0.79 $\pm$ 0.09 0.80 $\pm$ 0.09 0.71 $\pm$ 0.09
Single slice (2-D)
- 2-layer CNN 0.83 $\pm$ 0.04 0.84 $\pm$ 0.03 0.82 $\pm$ 0.05 0.83 $\pm$ 0.04 0.83 $\pm$ 0.03
Unimoal - VGG19 0.85 $\pm$ 0.03 0.85 $\pm$ 0.03 0.85 $\pm$ 0.04 0.85 $\pm$ 0.03 0.87 $\pm$ 0.03
- ResNet18 0.87 $\pm$ 0.04 0.86 $\pm$ 0.04 0.86 $\pm$ 0.04 0.88 $\pm$ 0.05 0.87 $\pm$ 0.04
Three slices (3-D)
MHCA [31] - 3D-CNN 0.80 $\pm$ 0.03 0.74 $\pm$ 0.04 0.88 $\pm$ 0.04 0.90 $\pm$ 0.03 0.71 $\pm$ 0.04
DeAF [32] - 3D-CNN 0.87 $\pm$ 0.03 0.84 $\pm$ 0.06 0.91 $\pm$ 0.03 0.93 $\pm$ 0.02 0.80 $\pm$ 0.06
TriFormer [19] Transformer 3D-CNN 0.83 $\pm$ 0.03 0.79 $\pm$ 0.03 0.89 $\pm$ 0.03 0.91 $\pm$ 0.02 0.75 $\pm$ 0.03
Proposed Model (2-D)
CGMCL GCN+GCN ResNet18 0.87 $\pm$ 0.02 0.86 $\pm$ 0.04 0.89 $\pm$ 0.03 0.89 $\pm$ 0.03 0.86 $\pm$ 0.03
CGMCL GCN+GCN VGG19 0.84 $\pm$ 0.01 0.82 $\pm$ 0.02 0.85 $\pm$ 0.03 0.85 $\pm$ 0.02 0.82 $\pm$ 0.02
CGMCL GAT+GAT VGG19 0.83 $\pm$ 0.04 0.83 $\pm$ 0.04 0.82 $\pm$ 0.04 0.83 $\pm$ 0.04 0.82 $\pm$ 0.04
CGMCL GAT+GAT ResNet18 0.89 $\pm$ 0.01 0.88 $\pm$ 0.02 0.91 $\pm$ 0.03 0.91 $\pm$ 0.03 0.88 $\pm$ 0.02
TABLE IV: The comparison of the proposed CGMCL’s performance (mean $\pm$ std) between unimodal and multimodal approaches for PD subtypes: Normal vs Abnormal. The best performance results are highlighted in bold, and “_” indicates the second-best methods
Normal vs. Abnormal
Models Backbone Image Feature Extractor Clinical Features ACC SEN SPE PPV NPV
Logistic - - 0.80 $\pm$ 0.00 0.80 $\pm$ 0.01 0.81 $\pm$ 0.01 0.80 $\pm$ 0.01 0.81 $\pm$ 0.01
XGboost - - 0.80 $\pm$ 0.00 0.77 $\pm$ 0.00 0.80 $\pm$ 0.02 0.78 $\pm$ 0.01 0.79 $\pm$ 0.03
AdaBoost - - 0.79 $\pm$ 0.04 0.78 $\pm$ 0.06 0.79 $\pm$ 0.05 0.78 $\pm$ 0.03 0.79 $\pm$ 0.06
Single slice (2-D)
- 2-layer CNN 0.86 $\pm$ 0.02 0.86 $\pm$ 0.03 0.87 $\pm$ 0.03 0.87 $\pm$ 0.03 0.86 $\pm$ 0.03
Unimodal - VGG19 0.87 $\pm$ 0.09 0.86 $\pm$ 0.01 0.88 $\pm$ 0.01 0.88 $\pm$ 0.02 0.86 $\pm$ 0.01
- ResNet18 0.90 $\pm$ 0.02 0.89 $\pm$ 0.02 0.91 $\pm$ 0.03 0.91 $\pm$ 0.03 0.89 $\pm$ 0.03
Three slices (3-D)
MHCA [31] - 3D-CNN 0.87 $\pm$ 0.01 0.92 $\pm$ 0.03 0.82 $\pm$ 0.02 0.85 $\pm$ 0.01 0.90 $\pm$ 0.03
DeAF [32] - 3D-CNN 0.88 $\pm$ 0.01 0.87 $\pm$ 0.02 0.89 $\pm$ 0.01 0.90 $\pm$ 0.00 0.86 $\pm$ 0.02
TriFormer [19] Transformer 3D-CNN 0.83 $\pm$ 0.03 0.86 $\pm$ 0.03 0.80 $\pm$ 0.03 0.83 $\pm$ 0.03 0.84 $\pm$ 0.03
Proposed Model (2-D)
CGMCL GCN+GCN ResNet18 0.87 $\pm$ 0.01 0.86 $\pm$ 0.02 0.89 $\pm$ 0.02 0.89 $\pm$ 0.02 0.86 $\pm$ 0.02
CGMCL GCN+GCN VGG19 0.88 $\pm$ 0.01 0.87 $\pm$ 0.02 0.89 $\pm$ 0.02 0.89 $\pm$ 0.02 0.87 $\pm$ 0.02
CGMCL GAT+GAT VGG19 0.89 $\pm$ 0.01 0.89 $\pm$ 0.02 0.89 $\pm$ 0.01 0.90 $\pm$ 0.01 0.89 $\pm$ 0.02
CGMCL GAT+GAT ResNet18 0.90 $\pm$ 0.01 0.90 $\pm$ 0.02 0.91 $\pm$ 0.02 0.91 $\pm$ 0.02 0.89 $\pm$ 0.02
TABLE V: Comparison of the accuracy of multi-class classification performance between unimodal and proposed CGMCL on melanoma dataset. The best performance results are highlighted in bold, and “_” indicates the second-best methods
Model Backbone Image Non-image BWV DaG PIG PN RS STR VS DIAG
Baseline
Logistic Model - 0.83$\pm$0.00 0.58$\pm$0.20 0.51$\pm$0.15 0.68$\pm$0.10 0.77$\pm$0.00 0.73$\pm$0.10 0.83$\pm$0.03 0.75$\pm$0.19
Xgboost - 0.82$\pm$0.00 0.49$\pm$0.14 0.68$\pm$0.12 0.67$\pm$0.09 0.76$\pm$0.00 0.74$\pm$0.10 0.84$\pm$0.03 0.74$\pm$0.19
AdaBoost - 0.83$\pm$0.03 0.54$\pm$0.09 0.64$\pm$0.04 0.60$\pm$0.03 0.77$\pm$0.02 0.68$\pm$0.04 0.84$\pm$0.01 0.66$\pm$0.03
Single image (2-D)
ResNet18 CNN 0.84$\pm$0.01 0.48$\pm$0.14 0.59$\pm$0.11 0.65$\pm$0.09 0.74$\pm$0.01 0.64$\pm$0.09 0.82$\pm$0.03 0.75$\pm$0.20
2-layer CNN 0.80$\pm$0.01 0.43$\pm$0.12 0.60$\pm$0.12 0.56$\pm$0.08 0.71$\pm$0.01 0.69$\pm$0.1 0.80$\pm$0.03 0.72$\pm$0.20
7-point [28] CNN 0.85 0.60 0.63 0.69 0.77 0.74 0.82 0.73
HcCNN [33] CNN 0.87 0.66 0.69 0.71 0.81 0.72 0.85 0.74
AMFAM [34] GAN 0.88 0.64 0.71 0.71 0.81 0.75 0.83 0.75
FusionM4Net [35] CNN 0.89$\pm$0.00 0.66$\pm$0.02 0.72$\pm$0.01 0.69$\pm$0.01 0.81$\pm$0.01 0.76$\pm$0.01 0.82$\pm$0.01 0.76$\pm$0.01
Proposed Model (2-D)
GCN+GCN 2-layer CNN 0.81$\pm$0.01 0.60$\pm$0.06 0.73$\pm$0.08 0.74$\pm$0.20 0.73$\pm$0.01 0.80$\pm$0.04 0.81$\pm$0.03 0.92$\pm$0.02
GAT+GAT 0.79$\pm$0.02 0.46$\pm$0.13 0.60$\pm$0.12 0.69$\pm$0.20 0.71$\pm$0.03 0.69$\pm$0.09 0.80$\pm$0.03 0.93$\pm$0.01
GCN+GCN ResNet18 0.87$\pm$0.01 0.65$\pm$0.01 0.76$\pm$0.03 0.75$\pm$0.05 0.78$\pm$0.02 0.76$\pm$0.02 0.86$\pm$0.02 0.95$\pm$0.01
GAT+GAT 0.87$\pm$0.01 0.67$\pm$0.03 0.74$\pm$0.03 0.75$\pm$0.03 0.78$\pm$0.02 0.77$\pm$0.02 0.87$\pm$0.01 0.94$\pm$0.01
Denotes the average of accuracy.
6 Ablation Study
In this section, we analyzed the impact of various parameters on CGMCL, including the classification performance with different values of the contrastive loss parameter $\beta$ across multiple backbone combinations and the parameter $K$ used for constructing the multi-modality neighborhoods graph, as shown in Supplementary Fig. S1.
6.1 Parameters of $\beta$ in the Objective Function
From a two-modality fusion perspective, we explore the influence of graph-based contrastive loss on feature representations in two domains. In Eq. 22, we introduce a parameterized $\beta$ as a weighting coefficient for $\mathcal{L}_{m}+\mathcal{L}_{f}$ and $\mathcal{L}_{contrastive}$. As Fig. 3, various CNN-based models serve as backbones, integrated with either GCN or GAT for graph-structured feature learning. Our results indicate that the combinations of ResNet18 with GAT or GCN achieve an AUC nearing 0.90 across different $\beta$ values in the MA vs. Abnormal and Normal vs. Abnormal classifications. However, in the most challenging task, Normal vs. MA, ResNet18 combined with GCN achieves a notable AUC of 0.80 at a $\beta$ value of 0.65.
6.2 Multimodal Features Alignment in KL Divergence Converge
The primary challenge in multimodal disease classification is ensuring proper alignment of features across different modalities during the fusion process. We evaluated the two learned representations $\hat{Z}^{I}$ and $\hat{Z}^{C}$ by converting them into probability matrices using the sigmoid function $\sigma(\cdot)$ and then calculating their Kullback-Leibler (KL) divergence, expressed as $\mathrm{KL}(\sigma(\hat{Z}^{I})\mathbin{\|}\sigma(\hat{Z}^{C}))$. The results are illustrated in Fig. 3 (b). A key observation is that the KL divergence across the three PD subtypes shows remarkable fluctuations after the first 50 epochs. This is especially apparent when comparing the Normal vs. MA subtypes, which require additional epochs to reach convergence. Melanoma showed clear convergence in terms of KL divergence. At the same time, PN and STR experienced a sharper decline in KL divergence during the first 50 epochs, indicating that the model can quickly differentiate these pathological markers. After 100 epochs, BWV exhibited more substantial changes in divergence compared to PN and STR. However, PN and STR demonstrated more stable convergence overall. PD maintained relatively low and stable KL divergence values, ranging from 0.4 to 0.7. In contrast, Melanoma initially had higher KL divergence values (near 1.2) but rapidly decreased and eventually stabilized between 0.4 and 0.6. These findings highlight how the multimodal model effectively captures distinct distributions among categories in these two domains and how CGMCL enhances overall training convergence.
TABLE VI: Evaluating classification performance and ablation study of module components in CGMCL
Model Backbone w/o Concatenate w/o IMFES CGMCL
Parkinson (Normal vs MA)
GCN+GCN 2-layer CNN $0.58\pm 0.07$ $0.53\pm 0.05$ 0.66 $\pm$ 0.03
GAT+GAT 2-layer CNN $0.62\pm 0.06$ $0.56\pm 0.08$ 0.66 $\pm$ 0.04
GCN+GCN ResNet18 $0.69\pm 0.04$ $0.71\pm 0.03$ 0.72 $\pm$ 0.05
GAT+GAT ResNet18 $0.73\pm 0.04$ $0.73\pm 0.04$ 0.77 $\pm$ 0.02
Parkinson (MA vs. Abnormal)
GCN+GCN 2-layer CNN $0.76\pm 0.12$ $0.63\pm 0.16$ 0.82 $\pm$ 0.01
GAT+GAT 2-layer CNN $0.80\pm 0.04$ $0.71\pm 0.13$ 0.85 $\pm$ 0.02
GCN+GCN ResNet18 $0.74\pm 0.13$ $0.82\pm 0.05$ 0.86 $\pm$ 0.02
GAT+GAT ResNet18 0.88 $\pm$ 0.02 $0.83\pm 0.03$ 0.85 $\pm$ 0.03
Parkinson (Normal vs. Abnormal)
GCN+GCN 2-layer CNN $0.84\pm 0.03$ $0.84\pm 0.02$ 0.86 $\pm$ 0.01
GAT+GAT 2-layer CNN $0.86\pm 0.02$ $0.77\pm 0.11$ 0.86 $\pm$ 0.02
GCN+GCN ResNet18 $0.79\pm 0.11$ $0.88\pm 0.02$ 0.89 $\pm$ 0.01
GAT+GAT ResNet18 $0.88\pm 0.02$ $0.87\pm 0.01$ 0.89 $\pm$ 0.01
Figure 4: The AUC and ACC performance of ablation experiments on various loss function combinations across the three PD subtypes.
6.3 Module Ablation
To validate CGMCL in module ablation for multimodal classification performance in latent space, we compare (1) w/o Concatenate, which excludes the concatenation from Eqs. 9, and  10, (2) w/o IMFES, which removes the IMFES module from Eqs. 11, and  12 under different backbones with the entire module. As shown in Table V, the experimental results indicate that the GAT+GAT combination with a ResNet18 backbone yields a slight AUC increase for the more distinct Normal vs. Abnormal subtypes, compared to models without concatenation or IMFES. Notably, CGMCL demonstrates a 5.5% improvement in the Normal vs. MA classification. Furthermore, it can inferred that utilizing various multi-level concatenation layers and weighted inter-modality distributions effectively enhances the extraction of critical feature information.
Figure 5: Visualization of CGMCL predictions incorporating six indicators from twelve DaTQUANT parameters.
6.4 Objective Function Ablation
We propose that a well-designed contrastive loss within the cross-graph modal objective function can effectively integrate feature spaces during the overall model optimization process. To evaluate CGMCL’s multimodal fusion performance, we tested various combinations of loss functions. These include the binary loss for single graph module fusion ($\mathcal{L}_{I}$ and $\mathcal{L}_{C}$), the addition of contrastive loss ($\mathcal{L}_{contrastive}$), and the comprehensive weighted loss CGMCL ($\mathcal{L}_{CGMCL}$). The experimental results in Fig. 4 show that $\mathcal{L}_{CGMCL}$ outperforms the combination of $\mathcal{L}_{I}$ + $\mathcal{L}_{C}$ by 2.1% in AUC and 2.2% in ACC for the MA vs. Abnormal classification task. Similarly, for the Normal vs. Abnormal task, $\mathcal{L}_{CGMCL}$ shows improvements of 1.1% in both AUC and ACC. These results indicate that the similarity between the two modalities for these subtypes is nearing the model’s convergence range. More importantly, the ability to distinguish early-stage PD is critical. In the Normal vs. MA classification task, $\mathcal{L}_{CGMCL}$ surpasses the next best-performing unweighted model $\mathcal{L}_{CGMCL}$ by 5.1% in AUC and 4.3% in ACC. This substantial improvement highlights CGMCL’s superior capability in capturing latent patterns characteristic of early-stage PD within the fusion space.
Figure 6: t-SNE visualization of CGMCL low-dimensional embeddings on two multimodal datasets.
Figure 7: Age-related regression estimates of CGMCL predictions for three PD subtypes.
7 Visualization of Multimodal Representation
7.1 Identification of PD subtype trajectories
For multimodal feature fusion, we present visual insights into the embeddings from the cross-graph modal fusion of patient images and meta-features generated by CGMCL. The final layer outputs of CGMCL serve as the two-dimensional embeddings for t-SNE visualization as shown in Fig. 5 and Fig. 6. We then project the corresponding images onto the low-dimensional data points. Fig. 6 (A) illustrates the low-dimensional embedding trajectories based on three severity classifications. The results show that CGMCL distinguishes between normal and abnormal cases in this reduced space. Our primary focus is the Normal vs. MA distinction, which is critical for early prevention. Our low-dimensional embedding demonstrates improved precision in differentiating subtle abnormalities from normal cases, addressing potential errors in manual classification.
7.2 Multi-Class Trajectories in Melanoma Embedding
For multi-class melanoma classification, the seven-point criteria categorize skin lesion characteristics into groups based on similarity, with physicians assigning corresponding scores. Fig. 6 (B) illustrates how CGMCL generates low-dimensional representations based on similar features for PIG, PN, and STR, which correspond to feature labels (ABS, RIG, IR), (ABS, TYP, ATP), and (ABS, REG, IR), respectively. In Fig. 6, for PN, the feature score of ATP is 2 (e.g., atypical pigment network, blue-whitish veil, and irregular vascular structures). It is evident that ATP has a more distinct and severe scoring, showing clear separation from ABS and TYP. Some ATP positions are located within TYP, indicating that certain TYP cases may progress toward more malignant features. Compared to low-dimensional representations in previous studies, our CGMCL model integrates more precise multimodal information into feature representations and provides diagnostic insights, addressing the lack of interpretability often associated with meta-features in existing research work.
Figure 8: Visualization of common similarity matrices for modality clustering results between patients i and j: Panel (a) shows the intrinsic similarity matrices for three PD patients. At the same time, panel (b) displays the similarity matrices of CGMCL clustering for different imbalanced classes in melanoma.
(a) Normal vs. MA
(b) MA vs. Abnormal
(c) Normal vs. Abnormal
Figure 9: The radar chart of explainable modality importance scores for quantitative analysis of overall striatal (R-L) indicators across different PD subtypes.
8 Discussion
8.1 Clustering Results
In this section, we assess the integration of diverse domains by evaluating CGMCL’s clustering capabilities for three subtypes of PD in similarity representations across various patient modalities. We visualize the patient similarity matrix derived from Eq. 13 for PD and melanoma in Fig. 8. Notably, CGMCL achieves near-perfect clustering between Normal and Abnormal cases in Fig. 8 (a). However, the clustering boundaries become less distinct when differentiating between Normal vs. MA and MA vs. Abnormal (e.g., prodromal and moderate abnormal PD). The broader clustering boundary range for Normal vs. MA in the prodromal stage indicates that CGMCL can, to some extent, distinguish prodromal PD patients. In the melanoma clustering analysis, as shown in Fig. 8 (b), the classification of different classes reveals three distinct community clusters in the similarity matrices of DaG and PIG across all categories. This demonstrates CGMCL’s ability to differentiate feature characteristics in multimodal melanoma, even when dealing with samples that have fewer classes.
8.2 Neurodegenerative Subtype Analysis in PD
Neurodegenerative conditions have long posed a predictive challenge in PD. Our proposed CGMCL aims to predict better and analyze the age-related effects on neurodegenerative PD subtypes. The scatter plot and regression estimates in Fig. 7 (left) reveal a weak but statistically significant correlation ($p<0.05$) between age and the model’s predictions for distinguishing Normal from MA. A subtle upward trend indicates that the model tends to classify older individuals as MA with a slightly higher probability. In contrast, the analysis of MA vs. Abnormal in Fig. 7 (middle) shows that age accounts for approximately 33.1% of the variance in model predictions, a finding that is highly statistically significant ($p<0.001$). This result indicates that age decreases the likelihood of being predicted as fully abnormal rather than mildly abnormal. Furthermore, Fig. 7 (right) demonstrates that the model’s predictions between Normal and Abnormal subtypes are not statistically significant ($p>0.05$), indicating that age is not a key factor in distinguishing these two categories. Although there is a slight increase in the probability of being classified as Abnormal with advancing age, the likelihood of being categorized as fully abnormal decreases with age. Based on these observations, the impact of age on differentiating abnormal cases appears minimal in late-stage PD, both in terms of manual interpretation and imaging features.
8.3 Explainable Modality Diagnosis
Traditional SPECT imaging for PD relies on manual interpretation and diagnosis by nuclear medicine physicians. In comparison to commonly used Grad-CAM methods (e.g., CNN visualization) for diagnosis [36], we developed a quantifiable diagnostic score that accurately predicts the magnitude of meta-features for patients at different stages. This score integrates our model’s predicted modality contribution with twelve meta-features, utilizing a scaled modality mechanism to calculate specific contribution scores [15]. In Fig. 9, we randomly selected patient samples from each subtype for analysis. We calculated the modality weight in CGMCL for the twelve meta-features, using importance scores given by $Score=X^{C}\odot\mathrm{IMFES}(\mathcal{H}^{C})$. Observing the Normal vs. MA comparison in Fig. 9 (a), patients beginning to exhibit MA characteristics (third image in Fig. 9 (a)) show a notably asymmetric geometric pattern in the striatal regions: PC-R, PC-L, and PP-L. An intriguing observation emerges when comparing the radar charts of normal patients with those of MA and Abnormal patients. Both groups with anomalies display substantially irregular radar chart shapes. This irregularity is particularly pronounced in the Abnormal group, where the overall area covered by the radar chart is smaller than in normal cases. This phenomenon is attributed to early-stage PD in patients, characterized by evident striatal atrophy in both hemispheres.
9 Conclusion
This study introduces a novel multimodal fusion framework that leverages cross-graph modal and integrates concatenated multi-level inner feature maps. Our approach effectively combines medical imaging data with clinical features, enhancing multimodal fusion. Additionally, we employ contrastive learning within the common latent space of fused same-modality data to improve the model’s classification accuracy for various subtypes across two multimodal datasets. Regarding clinical multimodal interpretability, our proposed CGMCL differs from existing methods by incorporating non-image features, particularly enabling quantitative clinical interpretation of SPECT imaging in early-stage PD. Furthermore, CGMCL shows potential for broader application in diverse multimodal disease studies.